Chattanooga is constructing a park at in investment of $1,000,000. Annual maintenance cost are expected to be $10,000 for the first 30 years and $20,000 thereafter for an indefinite period of time. With an interest rate of 6%, what is the capitalized cost of this project? Ans. $xxxxxxx
Answer #1

ANSWER:

capitalized cost of initial investment + capitalized cost of $10,000 adjustment for first 30 years will be

cc1 = -$1,000,000 + 10,000(p/a,6%,30) = -$1,000,000 + 10,000 * 13.765 = -$1,000,000 + $137,650 = -$862,350

capitalized of $20,000 maintenance forever = maintenance cost / interest rate = -$20,000 / 6% = -$333,333.33

total capitalized cost of the project = cc1 + cc2 = -$862,350 + (-$333,333.33) = -$862,350 - $333,333.33 = -$1,195,683.33